### v1

# Beyond Pixels: Efficient Dataset Distillation via Sparse Gaussian Representation

Abstract
Dataset distillation has emerged as a promising paradigm that synthesizes compact, informative datasets capable of retaining the knowledge of large-scale counterparts, thereby addressing the substantial computational and storage burdens of modern model training. Conventional approaches typically rely on dense pixel-level representations, which introduce redundancy and are difficult to scale up. In this work, we propose GSDD, a novel and efficient sparse representation for dataset distillation based on 2D Gaussians. Instead of representing all pixels equally, GSDD encodes critical discriminative information in a distilled image using only a small number of Gaussian primitives. This sparse representation could improve dataset diversity under the same storage budget, enhancing coverage of difficult samples and boosting distillation performance. To ensure both efficiency and scalability, we adapt CUDA-based splatting operators for parallel inference and training, enabling high-quality rendering with minimal computational and memory overhead. Our method is simple yet effective, broadly applicable to different distillation pipelines, and highly scalable. Experiments show that GSDD achieves state-of-the-art performance on CIFAR-10, CIFAR-100, and ImageNet subsets, while remaining highly efficient encoding and decoding cost. Our code is available at https://github.com/j-cyoung/GSDatasetDistillation.
1 introduction
Guided by the principles of scaling laws Kaplan et al. (2020), deep learning has advanced along the trajectory of larger models, larger datasets, and longer training, which has significantly pushed the performance boundaries of modern models. However, this paradigm also incurs enormous demands on computation, storage, and communication resources, creating a barrier that hinders both broad adoption and sustainable development. Consequently, a pivotal question in the field is how to achieve strong performance under limited computational and data resources.
Against this backdrop, dataset distillation offers a promising solution. Its goal is to distill the essential knowledge of massive datasets into a compact set of synthetic samples, enabling models trained on these small sets to approximate the performance of those trained on the full dataset. By adopting dataset distillation, the costs of model training, data storage, and data transmission can be substantially reduced. Moreover, it opens up a wide range of applications, including efficient data replay for continual learning Yang et al. (2023); Gu et al. (2024), critical data transfer in federated learning Huang et al. (2024a), and privacy preservation Dong et al. (2022).
Dataset distillation optimizes the synthetic data itself, with the objective of minimizing the training-aware information gap between the synthetic and original datasets. Broadly, dataset distillation can be decomposed into two complementary modules: distillation algorithms and data parameterization. The former defines how training-aware key information, such as training gradients Zhao et al. (2021), model parameters Cazenavette et al. (2022), or distributional statistics Zhao & Bilen (2023), is extracted, while the latter determines the representation format of the synthetic data. Although extensive research has focused on innovating distillation algorithms, exploration of the representation space has been relatively limited, with most approaches directly optimizing in the pixel space. Conventional methods typically adopt a naïve pixel grid as the parameterization scheme. Such dense representations fail to capture the relative importance of pixels and lead to substantial storage redundancy. Decoder-based parameterizations, in contrast, can leverage structural priors to enhance image realism and share common information, but they inevitably restrict the optimization space and introduce additional computational and memory overhead. More recently, approaches based on implicit neural representations (INRs), such as DDiF Shin et al. (2025), have substantially reduced per-image storage and thereby improved the diversity of distilled datasets. However, their inherent per-pixel query decoding mechanism incurs prohibitive computational costs, particularly in high-resolution or large-batch scenarios, creating a severe performance bottleneck.
In this paper, we propose a new paradigm for dataset distillation, termed Gaussian Splatting Dataset Distillation (GSDD). Instead of relying on dense pixel grids, GSDD represents each distilled image with a sparse set of 2D Gaussians. Each Gaussian explicitly encodes training-aware image features that span multiple pixels, and optimizes only a few parameters such as position and shape, thereby replacing large-scale pixel-level optimization. This sparse representation significantly reduces the storage cost per image, which in turn increases the diversity of the distilled dataset and improves coverage across samples of varying difficulty. To further ensure the scalability of GSDD, we adapt a highly parallelized differentiable rasterizer that enables instantaneous high-quality rendering from parameters to images.
Our contributions can be summarized as follows:
  • We propose a novel dataset parameterization framework based on 2D Gaussian representations. This is the first work that introduces sparse Gaussian representations into dataset distillation. The method is both simple and effective, and it can enhance the performance of a wide range of distillation algorithms.
  • We provide a complete design of the Gaussian representation, including parameterization, optimization objectives, and an efficient batch-parallel Gaussian rasterization operator. This design ensures fast convergence, high distillation quality, and stable optimization.
  • Extensive experiments demonstrate that our method achieves state-of-the-art performance on CIFAR-10, CIFAR-100, and multiple ImageNet subsets, while offering faster computation and lower memory consumption compared with prior approaches.
2 Background and Motivation
2.1 Related Work
Dataset Distillation Algorithms
Modern deep learning models typically rely on large-scale training datasets, whereas the goal of dataset distillation is to replace the dataset with a much smaller set of high-quality training samplesWang et al. (2018). Unlike data selection, dataset distillation explicitly optimizes the synthetic data to achieve superior performance.
Formally, let the original dataset be denoted as
$\mathcal{T}=\{(t_{i},y_{i})\mid i=1,\cdots,N_{\mathcal{T}}\}\subseteq\mathcal{D},$
where $\mathcal{D}$ represents the underlying data distribution. The distilled dataset (also referred to as the synthetic dataset) is defined as
$\mathcal{S}=\{(s_{j},y_{j})\mid j=1,\cdots,N_{\mathcal{S}}\},$
where $t_{i}$ and $s_{j}$ denote original and distilled images, and $y_{i}$ and $y_{j}$ are their corresponding labels. Since this work adopts one-hot label encoding, without loss of generality we denote the distilled dataset as
$\mathcal{S}=\{s_{j}\mid j=1,\cdots,N_{\mathcal{S}}\}.$
The storage footprint of $\mathcal{S}$ is typically required to be far smaller than that of $\mathcal{T}$. In dataset distillation, the synthetic dataset is treated as a set of learnable objects, where the pixel values of distilled images $s_{j}$ are regarded as optimization parameters. The ultimate goal is to train a model $\theta_{\mathcal{S}}$ on the compact distilled dataset such that its performance closely matches that of a model $\theta_{\mathcal{T}}$ trained on the full dataset Wang et al. (2018):
$\mathcal{S}=\arg\min_{\mathcal{S}}\ \big|l(\theta_{\mathcal{S}},\mathcal{D})-l(\theta_{\mathcal{T}},\mathcal{D})\big|,\quad\theta_{\mathcal{S}}=\arg\min_{\theta}l(\theta,\mathcal{S}),\quad\theta_{\mathcal{T}}=\arg\min_{\theta}l(\theta,\mathcal{T}).$
This optimization objective can be approached via meta-learning, but such methods often consume substantial GPU memory Feng et al. (2024) and yield suboptimal performance due to weak supervision signals. To address these challenges, many alternative distillation objectives have been proposed. The core idea is to extract and align critical training-aware information between $\mathcal{S}$ and $\mathcal{T}$. Mainstream approaches include trajectory matching(TM) Cazenavette et al. (2022); Du et al. (2023); Liu et al. (2025); Guo et al. (2024); Cui et al. (2023), which constrains the distance between model parameters trained on the two datasets; distribution matching(DM) Zhao & Bilen (2023); Zhao et al. (2023); Liu et al. (2023); Wei et al. (2024); Wang et al. (2025b), which aligns feature distributions; and gradient matching(DC) Zhao et al. (2021); Lee et al. (2022); DU et al. (2023), which constrains the gradients derived from the two datasets. Despite the drastic reduction in storage requirements, distilled datasets often achieve competitive training performance, and the resulting synthetic images can diverge significantly from the originals. Consequently, dataset distillation has found applications in continual learning Yang et al. (2023); Gu et al. (2024), privacy preservation Dong et al. (2022); Chung et al. (2024); Zheng et al. (2025), federated learning Huang et al. (2024a); Jia et al. (2025); Yan et al. (2025), and neural architecture search Such et al. (2020).
Parameterization of Dataset Distillation
The parameterization of the distilled dataset $\mathcal{S}$ critically influences both performance and efficiency. Early methods directly optimized RGB pixels, but this dense representation wastes storage by ignoring spatial redundancy. To address this, decoder-based approaches Liu et al. (2022); Liu & Wang (2023); Cazenavette et al. (2023) learn latent codes with a shared decoder, which reduces redundancy but limits flexibility in optimization and adds computation costs. Dictionary-based methods Deng & Russakovsky (2022); Wei et al. (2023) instead combine basis images with index matrices to support domain adaptation and generalization, though they still store pixel-level bases and are restricted by linear combinations.
Recent studies have introduced compact low-level representations to reduce image redundancy and increase distilled samples within a fixed storage budget, such as downscaling Kim et al. (2022), frequency-domain transforms Shin et al. (2023), and implicit neural representations (INRs) Shin et al. (2025). INR-based methods show strong performance by encoding each image as a neural field with fewer parameters, but their efficiency is limited since decoding requires per-pixel RGB queries, causing heavy overhead for high-resolution and large-batch training.
Gaussian Splatting
Gaussian splatting Kerbl et al. (2023) represents 3D scenes using Gaussian ellipsoids, offering higher rendering efficiency and explicit geometric control compared to NeRF-based representations Mildenhall et al. (2021). Originally developed for novel view synthesis, it initializes a set of Gaussian primitives within a 3D scene, which are then rendered from multiple viewpoints and optimized to match the corresponding ground-truth images, ultimately enabling arbitrary-view synthesis. In recent years, Gaussian splatting has been extended to a wide range of domains, including digital humans Qian et al. (2024); Zielonka et al. (2025), autonomous driving Zhou et al. (2024); Huang et al. (2024b), dynamic scene modeling Wu et al. (2024); Huang et al. (2024c), and 3D editing Chen et al. (2024). More recently, Gaussian representations have also been applied to 2D vision tasks such as image super-resolution Hu et al. (2024); Chen et al. (2025), image representation Zhu et al. (2025); Weiss & Bradley (2024); Zhang et al. (2025), and video representation Wang et al. (2025a), where Gaussian primitives are used to efficiently encode high-fideliy image of videos with fine-grained textures.
This work is the first to introduce Gaussian Splatting as a parameterization for dataset distillation, which leverages Gaussian representations as a simple yet efficient parameterization for distilled datasets. Unlike prior approaches that rely on network structures or complex operations, GSDD employs a CUDA-based differentiable rasterization algorithm to render sparse Gaussian primitives. As an explicit representation, it offers high flexibility in optimization, substantially increases dataset diversity under the same storage budget, and significantly accelerates both encoding and decoding. Rather than emphasizing fine-grained image fidelity, our method parameterizes entire distilled datasets with Gaussian representations and directly optimizes them for model training performance.
2.2 Motivation
Gaussian representations enable faster decoding and rendering, especially compared to the state-of-the-art DDiF Shin et al. (2025) method. DDiF relies on implicit neural representations using a SIREN network Sitzmann et al. (2020) to model distilled images, where each pixel’s RGB value must be queried individually, as illustrated in Figure 1a. This approach incurs substantial time and memory overhead as the number and resolution of distilled images increase. In contrast, GSDD implements a custom CUDA operator for batched rendering of Gaussians. This allows all Gaussians to be efficiently rendered into distilled images in a single pass. As shown in Figure 1b, this design significantly reduces both inference and optimization time and memory usage, thereby improving the scalability of dataset distillation methods.
Figure 1: Comparison of Different Distilled Image Representation
Gaussian representations enable compact modeling of individual distilled images with very few parameters. Traditional dataset distillation methods operate in the pixel space, where each pixel is assigned the same storage budget and optimized equally. This uniform treatment introduces unnecessary storage and computation costs. In contrast, GSDD can use a single Gaussian to represent a larger spatial area, covering multiple pixels simultaneously. This not only provides a more efficient image representation but also enhances optimization efficiency. As shown in Figure 1c, pixel-based methods optimize each pixel independently, which inevitably introduces redundancy and noise. In comparison, Gaussian-based optimization aggregates pixel-level gradients within the support region of each Gaussian, resulting in more robust updates. Furthermore, each Gaussian can perform localized optimization by adjusting a small set of parameters such as position, color, and shape. This region-level update strategy improves both efficiency and effectiveness during training.
3 methodology
Figure 2: Overview of the proposed framework. Each Gaussian is parameterized by a total of 9 floating-point values. A single distilled image is represented by a set of Gaussians. During training, we first quantize the parameters to bf16 precision to obtain quantized Gaussian primitives. These are then rendered into distilled images using a customized rasterizer. During the rasterization process, we concatenate all primitives from the distilled dataset and feed them into a single batched rasterization kernel. This design enables efficient rendering and facilitates a compact data structure, as the entire distilled dataset can be managed by initializing only one object instance. To further improve rendering quality when Gaussians are sparse, we incorporate prefiltering and SuperSampling-based Anti-Aliasing techniques. These enhancements enable more accurate estimation of RGB values at each pixel. The rendered distilled images are then aligned with the original images for information matching, and the resulting gradients are backpropagated to update the Gaussian parameters.
3.1 Parameterization of Dataset Distillation By Gaussian Splatting
We adopt a 2D Gaussian mixture model to parameterize distilled images. Each distilled image $s_{j}$ is represented as a set of $M$ Gaussian components, denoted by
$G_{j}=\{g_{k}\mid k=1,\cdots,M\}.$ (1)
Specifically, for a pixel located at coordinates $(x,y)$, the contribution from the $k$-th Gaussian component $g_{k}$ is defined by the following unnormalized Gaussian function:
$g_{k}(x,y;\mu_{k},\Sigma_{k})=\exp\!\left(-\tfrac{1}{2}(\mathbf{x}-\mu_{k})^{T}\Sigma_{k}^{-1}(\mathbf{x}-\mu_{k})\right),$ (2)
where $\mathbf{x}=[x,y]^{T}$ denotes the pixel center and $\mu_{k}=[u_{k},v_{k}]^{T}$ is the Gaussian mean. The covariance matrix $\Sigma_{k}$ determines the size, shape, and orientation of the Gaussian. To ensure that $\Sigma_{k}$ remains positive semi-definite during optimization, we parameterize it via a stable Cholesky decomposition, i.e., $\Sigma_{k}=L_{k}L_{k}^{T}$, where the lower-triangular matrix $L_{k}$ is given by
$L_{k}=\begin{bmatrix}l_{11}^{k}&0\\ l_{21}^{k}&l_{22}^{k}\end{bmatrix}$. (3)
Each Gaussian $g_{k}$ is associated with a color vector $\mathbf{c}_{k}\in\mathbb{R}^{3}$ and an opacity $\alpha_{k}$. The final color of pixel $(x,y)$ is then synthesized by aggregating the contributions of all Gaussians:
$\mathbf{c}(x,y)=\sum_{k=1}^{M}\alpha_{k}\cdot g_{k}(x,y;u_{k},v_{k},\Sigma_{k})\cdot\mathbf{c}_{k}.$ (4)
Thus, each Gaussian component can be fully described by a 9-dimensional vector $p_{k}=(u_{k},v_{k},l_{11}^{k},l_{21}^{k},l_{22}^{k},\mathbf{c}_{k},\alpha_{k})\in\mathbb{R}^{9}.$ Accordingly, the parameter set of an entire distilled image $s_{j}$ is given by $\mathbf{p}^{j}=\{p_{k}^{j}\}_{k=1}^{M}$, and the parameterization of the whole distilled dataset $\mathcal{S}$ is $\mathcal{P}_{\mathcal{S}}=\{\mathbf{p}^{j}\}_{j=1}^{N_{\mathcal{S}}}.$ Based on this parameterization, the rendering process of a distilled image can be expressed as a differentiable synthesis function
$s_{j}=R(\mathbf{p}^{j};\mathcal{C}),$ (5)
which maps the parameters $\mathbf{p}^{j}$ onto a predefined pixel coordinate grid $\mathcal{C}$. Here, $\mathcal{C}$ denotes the set of all pixel coordinates, i.e.,
$\mathcal{C}=\{x\mid x\in 1/W,\cdots,(W-1)/W\}\times\{y\mid y\in 1/H,\cdots,(H-1)/H\}.$ (6)
This Gaussian-based parameterization is orthogonal to the choice of dataset distillation algorithm. Once a differentiable synthesis function is defined, it can be seamlessly integrated into any distillation framework. The optimization objective is to find the optimal distilled parameters $\text{Param}(\mathcal{S})^{*}$ that minimize the distillation loss $L_{\text{distill}}$:
$\text{Param}(\mathcal{S})^{*}=\arg\min_{\text{Param}(\mathcal{S})}L_{\text{distill}}(\mathcal{S},\mathcal{T}),$ (7)
where $L_{\text{distill}}$ can correspond to any distillation objective.
To accelerate convergence and provide a good initialization for the optimization process, we pre-initialize the parameters of the distilled dataset. Specifically, we randomly sample a small subset of real images $\mathcal{S}_{0},|\mathcal{S}_{0}|=N_{\mathcal{S}}$ from the original dataset $\mathcal{T}$, and minimize the mean squared error (MSE) between the synthesized images and these real samples. The initialization objective is thus:
$\text{Param}(\mathcal{S})_{0}=\arg\min_{\text{Param}(\mathcal{S})}L_{\text{MSE}}(\mathcal{S},\mathcal{S}_{0}).$ (8)
3.2 Efficient Gaussian Splatting Design
The core of the above optimization lies in the differentiable renderer $R(\mathbf{p}^{j};\mathcal{C})$, which maps Gaussian primitives to distilled images with both high visual fidelity and computational scalability. To enable efficient processing of Gaussian representations for large-scale dataset distillation, we incorporate several critical designs into the rendering pipeline, including parallel rendering of multiple distilled images, anti-aliasing strategies, spatial constraints on Gaussian positions and quantization.
Existing open-source rasterization operators are primarily designed for 3D scene reconstruction or single-image tasks such as representation learning and super-resolution, and generally lack support for multi-image rendering. To fully leverage the parallelism of modern GPUs, we implement customized data structures and rendering kernels for both the forward and backward passes. Specifically, all Gaussian primitives across the entire distilled dataset are represented as a single contiguous 1D vector, and a globally unique ID (GUID) system is used to assign threads that rasterize multiple images in parallel. Details are provided in Appendix A.3, and code will be open-sourced.
Some Gaussians may become highly anisotropic during optimization to capture high-frequency features such as edges, which necessitates anti-aliasing strategies. We adopt two complementary techniques: (i) analytic pre-filtering, which approximates each Gaussian’s integral over a pixel area by convolving it with a unit pixel box filter, resulting in a modified covariance
$\Sigma_{k}^{\prime}=\Sigma_{k}+\Sigma_{box},\quad\text{where}\quad\Sigma_{box}=\text{diag}\left(\tfrac{1}{12},\tfrac{1}{12}\right),$ (9)
introducing a minimum rendering variance that suppresses aliasing for extremely narrow Gaussians; and (ii) $\times2$ supersampling anti-aliasing (SSAA), which averages multiple samples within each pixel area to smooth boundaries and reduce high-frequency artifacts with minimal overhead.
During optimization, some Gaussian centers tend to drift outside the normalized coordinate space $1,1]^{2}$, where they receive no gradients and become unrecoverable, a phenomenon we term Gaussian Escape. To alleviate this, we apply a boundary regularization loss to slightly encourage Gaussian centers to remain within the viewable region:
$l_{\text{boundary}}=-\mathbb{E}_{k}\big[\log(1-\tilde{x}^{2})+\log(1-\tilde{y}^{2})\big].$ (10)
In practice, the positions of Gaussian primitives are defined within the range [1,1][-1,1], and we perform coordinate transformations inside the CUDA operator. At each iteration, we clip the Gaussian positions to ensure they remain within the valid bounds:
$\tilde{\mu}_{k}=\text{clip}(\mu_{k},-1+\epsilon,1-\epsilon).$ (11)
We observe that Gaussian parameters exhibit robustness to reduced numerical precision, as the rasterization process does not suffer from cumulative precision errors. Based on this, we store all Gaussian parameters in bfloat16 (bf16) precision. During training, parameters are maintained in fp32 to ensure accurate updates, while bf16 casting is applied during the forward pass to allow the model to adapt to quantization effects. After training, all parameters are quantized to bf16.
3.3 Analysis of the Gaussian Representation
Figure 3: (a) Distillation performance under different Gaussian pruning strategies as a function of the remaining Gaussian ratio; (b) Test accuracy across training epochs with different GPC (Gaussian Images Per Class) under the same storage budget; (c) Test accuracy of the distilled dataset on samples of varying difficulty under the same storage budget; (d) Relationship between prediction accuracy on samples of different difficulty and GPC under the same storage budget. For fair comparison, TM is initialized with real images and serves as a baseline that represents pixel-based distillation.
Sparsity improves optimization
A key advantage of using Gaussian representations for image distillation lies in their ability to efficiently capture and encode training-aware image features that span multiple pixels, using only a single or very few Gaussian primitives. As a result, this approach yields a more compact and efficient representation for each distilled image. We design a pruning experiment based on the following hypothesis: Gaussian components with larger spatial extent and higher opacity are more likely to carry critical training information. As shown in Figure 3(a), when these large and opaque Gaussians are pruned first, the model performance drops sharply as the pruning ratio increases. In contrast, removing smaller and more transparent Gaussians has less effect on the performance. The results indicate that large Gaussian components not only cover broader pixel regions but also contribute more significantly to the final performance of the distilled dataset. This further confirms that the proposed Gaussian representation captures a highly sparse and information-dense encoding of the distilled image.
Figure 4: (a) Loss landscape of GSDD; (b) Loss landscape of pixel-based representation.
The Gaussian representation enables efficient modeling of distilled images with a small number of parameters. Its core mechanism lies in the ability to apply structured and coherent modifications to an entire pixel region by adjusting the properties of a single Gaussian component, such as its position, shape, color, and opacity. During backpropagation, each Gaussian also naturally aggregates the gradients of all pixels it covers. This optimization paradigm, which operates at the level of geometric primitives, offers improved robustness and efficiency compared to traditional pixel-wise optimization. As shown by the convergence curves in Figure 3(b), GSDD achieves faster convergence and higher final performance than pixel-based methods. To further examine the underlying differences in optimization behavior, we visualize the loss landscapes of both approaches in Figure 4(a) and 4(b). GSDD exhibits a smoother loss surface with a lower minimum and a wide basin resembling that of convex optimization, which substantially reduces the difficulty of gradient descent. In contrast, the pixel-based method presents a highly rugged and noisy loss landscape, making the optimization process more susceptible to poor local minima.
Diversity improves scalability
The parameter efficiency of Gaussian representations allows more distilled images to be synthesized under a fixed storage budget, thereby improving coverage of and generalization to more challenging samples. Prior work has shown that models often rely on memorization to generalize to hard examples, a phenomenon that is difficult to address in conventional dataset distillation settings. To evaluate this advantage, we conduct the following experiment: we first construct a difficulty spectrum by counting the number of incorrect predictions made by 100 independently trained models for each test sample. We then fix the overall storage budget and compare the performance of student models trained on distilled datasets with different Gaussian images Per Class(GPC) values, measuring their accuracy across different difficulty ranges. As shown in Figures 3(c) and 3(d), the results reveal a clear trend: increasing the number of representable distilled images initially yields substantial performance gains on easy samples. With higher GPC values, the models further improve their accuracy on harder examples. These findings suggest that GSDD enhances downstream generalization by increasing the diversity of distilled images, thereby covering a broader range of the difficulty spectrum.
4 experiments
Table 1: Results on ImageNet-Subset, TM
Subset Nette Woof Fruit Yellow Meow Squawk
Original 87.4 67.0 63.9 84.4 66.7 87.5
Input sized TM (Vanilla) 51.4 29.7 28.8 47.5 33.3 41.0
FRePo 48.1 29.7
Static IDC 61.4 34.5 38.0 56.5 39.5 50.2
FreD 66.8 38.3 43.7 63.2 43.2 57.0
Parameterized HaBa 51.9 32.4 34.7 50.4 36.9 41.9
SPEED 66.9 38.0 43.4 62.6 43.6 60.9
NSD 68.6 35.2 39.8 61.0 45.2 52.9
Generative Prior GLaD 38.7 23.4 23.1 26.0 35.8
H-GLaD 45.4 28.3 25.6 29.6 39.7
Function DDiF 72.0 42.9 48.2 69.0 47.4 67.0
Primitive GSDD 76.4 46.4 51.2 72.4 53.8 73.2
Table 2: Results with DC/DM
DC ImageNet Subset ($\times128$) Avg
Nette Woof Fruit Yellow Meow Squawk
DC (Vanilla) 34.2 22.5 21.0 37.1 22.0 32.0 28.1
GLaD 35.4 22.3 20.7 22.6 33.8 27.0
H-GLaD 36.9 24.0 22.4 24.1 35.3 28.5
IDC 45.4 25.5 26.8 25.3 34.6 31.5
FreD 49.1 26.1 30.0 28.7 39.7 34.7
DDiF 61.2 35.2 37.8 39.1 54.3 45.5
GSDD 70.2 42.4 47.8 64.0 43.2 69.4 56.2
DM
DM (Vanilla) 30.4 20.7 20.4 36.0 20.1 26.6 25.7
GLaD 32.2 21.2 21.8 22.3 27.6 25.0
H-GLaD 34.8 23.9 24.4 24.2 29.5 27.4
IDC 48.3 27.0 29.9 30.5 38.8 34.9
FreD 56.2 31.0 33.4 33.3 42.7 39.3
DDiF 69.2 42.0 45.3 45.8 64.6 53.4
GSDD 74.6 43.4 52.0 69.4 48.2 73.6 60.2
4.1 Experimental Results
Datasets and Baselines
We evaluate our method on standard datasets: CIFAR-10, CIFAR-100 (at $32\times 32$ resolution), and six ImageNet subsets (at $128\times 128$ resolution). The general pipeline of dataset distillation involves two stages: (1) generating a synthetic dataset, and (2) training a model from scratch on the synthetic data. The performance of the trained model on the real test set is used to assess the quality of the distilled dataset. We include a wide range of recent dataset distillation baselines, including TM Cazenavette et al. (2022), FRePo Zhou et al. (2022), IDC Kim et al. (2022), FreD Shin et al. (2023), HaBa Liu et al. (2022), RTP Deng & Russakovsky (2022), HMN Liu et al. (2022), SPEED Wei et al. (2023), NSD Yang et al. (2025), GLaD Cazenavette et al. (2023), H-GLaD Zhong et al. (2025), LD3M Moser et al. (2024) and DDiF Shin et al. (2025).
Experimental Details
We use TM Cazenavette et al. (2022), DM Zhao & Bilen (2023), and DC Zhao et al. (2021) as the underlying distillation algorithms in our experiments. All parameters of Gaussians are trained using the Adam optimizer, and the learning rate for Gaussian parameters is uniformly set to 0.001. For CIFAR datasets, we apply ZCA whitening as a standard preprocessing step. We fix the boundary loss weight to $\lambda_{\text{boundary}}=0.1$. Experiments are conducted on NVIDIA V100 and RTX 4090 GPUs. To initialize the Gaussian representation, we randomly sample real images from the original dataset and fit them with Gaussians before entering the standard distillation loop. We provide complete implementation details and hyperparameters in the Appendix A.1.
Table 3: Ablation
ImageNette ACC
w/o opacity 74.8
w/o boundary 75.1
w/o bf16 74.8
w/o antialias 76.4
all 76.4
CIFAR-10 ACC
w/o antialias 74.5
all 75.5
Storage Budget and Evaluation Protocol
We evaluate performance under varying storage budgets, specified in IPC (images per class). Since our method often utilizes multiple low-storage Gaussian images, we additionally report GPC (Gaussian Images per Class) in Appendix A.1. The number of Gaussians per image is computed as $\text{pts}=\frac{\text{res}\times\text{res}\times 3\times\text{IPC}\times 2}{\text{GPC}\times 9}$, where the factor of 2 accounts for our use of bf16 to store each Gaussian parameter, and the denominator 9 reflects the total number of parameters per Gaussian.
Performance Comparison
Through the integration of GSDD, our method consistently outperforms the pixel-based TM baseline and other dataset parameterization methods on six ImageNet subsets with IPC=1, as shown in Table 2. Extensive results on CIFAR-10 and CIFAR-100 with IPC values of 1, 10, and 50 are provided in Appendix A.2. When compared to the state-of-the-art method DDiF, our approach achieves consistent improvements, demonstrating highly competitive performance. These results highlight the effectiveness of GSDD in enhancing the quality of distilled datasets.
Table 4: Cross Architecture Results
Subset Nette Woof Fruit Yellow Meow Squawk
TM 22.0 14.8 17.1 22.3 16.2 25.5
IDC 27.9 19.5 23.9 28.0 19.8 29.9
FreD 36.2 23.7 23.6 31.2 19.1 37.4
GLaD 30.4 17.1 21.1 19.6 28.2
H-GLaD 30.8 17.4 21.5 20.1 28.8
LD3M 32.0 19.9 21.4 22.1 30.4
DDiF 59.3 34.1 39.3 51.1 33.8 54.0
GSDD 58.1 34.6 39.9 53.6 34.0 58.0
Universality to Matching Objectives
GSDD is designed to be compatible with a wide range of dataset distillation algorithms and can be directly integrated with them. We evaluate GSDD under two widely adopted loss paradigms: gradient matching(DC) and distribution matching(DM). Experimental results, as shown in Table 2, confirm that GSDD robustly enhances performance across different distillation objectives.
Cross-Architecture Performance
An essential property of dataset distillation is robustness across model architectures. The distilled data should retain high performance even when the evaluation model differs from the one used during distillation. To assess this, we synthesize data using a ConvNet and evaluate it on a variety of downstream architectures, including ResNet, VGG, and others. The results, summarized in Table 4, demonstrate that the increased diversity introduced by GSDD leads to substantial improvements in cross-architecture generalization.
4.2 More Analysis
Figure 5: Performance comparison between GSDD and DDiF under the same per-image storage budget (abbreviated as st (floats)). Top row: Forward and forward+backward execution time and memory usage across varying image resolutions (with fixed batch size = 32). Bottom row: Same metrics across varying batch sizes (with fixed resolution = 128). GSDD consistently achieves lower latency and memory consumption, especially under high-resolution and large-batch scenarios.
To investigate the contribution of each component to model performance, we conduct a series of ablation studies, with results summarized in Table 3. The experiments show that opacity modeling and bf16 mixed-precision training are two major factors, as removing either leads to a notable performance drop. The impact of anti-aliasing exhibits a dependence on the dataset and the complexity of the representation. Its effect becomes important when the number of Gaussians used to represent each image is limited. Specifically, for datasets like CIFAR-10, where storage budgets are constrained and each image is represented by a small number of Gaussians, disabling anti-aliasing introduces rendering artifacts that degrade performance. In contrast, for datasets like ImageNette with denser Gaussians, aliasing artifacts are largely mitigated, and the performance gains from anti-aliasing become relatively minor.
To comprehensively evaluate the efficiency of our approach, we conduct a quantitative comparison with the state-of-the-art DDiF method in terms of both decoding and training performance. As shown in Figure 5, we measure the inference time and memory consumption of both methods across varying per-image storage budgets, image resolutions, and batch sizes. The results demonstrate that GSDD consistently outperforms DDiF in both inference/training speed and memory efficiency. This advantage becomes especially pronounced when handling high-resolution inputs or large batch sizes, where GSDD achieves several-fold, or even an order-of-magnitude, reductions in both computation time and memory usage. These findings indicate that our method not only maintains strong image representation capabilities for distillation but also offers superior rendering speed and significantly lower memory footprint, thereby substantially improving the scalability of dataset distillation methods in practical deployment scenarios.
5 conclusion
This work introduces sparse Gaussian representations into the task of dataset distillation parameterization. We propose a complete framework for encoding distilled images using 2D Gaussian primitives, including the underlying mathematical formulation, a parallel differentiable renderer, and several tailored components to handle the sparsity of distilled datasets. Specifically, we incorporate anti-aliasing and spatial constraints to enhance the utility of Gaussians and improve overall distillation quality. Through extensive quantitative and qualitative experiments, we demonstrate that sparse Gaussian representations facilitate more effective optimization, offer strong scalability, and provide better coverage across samples of varying difficulty. As shown in our results, GSDD achieves highly competitive performance under various datasets and storage budgets, while also exhibiting superior computational efficiency compared to DDiF. The proposed method is simple, plug-and-play, and readily compatible with existing distillation algorithms, leaving ample room for future extensions. Future work may focus on scaling to larger datasets, extending Gaussian representations to video modalities, and exploring dynamic density control of Gaussian primitives to further enhance representation quality.
Appendix A Appendix
A.1 Experimental Details
A.1.1 Datasets
We conduct experiments on eight datasets in total, including CIFAR-10, CIFAR-100, and six ImageNet subsets: imagenette, imagewoof, imagefruit, imagemeow, imagesquawk, and imageyellow. CIFAR-10 consists of 60,000 images of resolution $32\times 32$ across 10 classes, with 50,000 training images and 10,000 test images. CIFAR-100 has the same resolution and number of images, but covers 100 classes. Each ImageNet subset contains images of resolution $128\times 128$ from 10 classes, with more than 10,000 images per subset.
A.1.2 Network Architectures
We adopt ConvNet architectures as the backbone networks. A Depth-$n$ ConvNet consists of $n$ blocks followed by a fully-connected layer, where each block is composed of a $3\times 3$ convolutional layer with 128 filters, instance normalization, a ReLU activation, and $2\times 2$ average pooling with stride 2. Specifically, we use ConvNetD3 for CIFAR-10 and CIFAR-100, and ConvNetD5 for the ImageNet subsets.
A.1.3 Image Initialization
For the initialization of synthetic images, all experiments adopt the mean squared error (MSE) loss and use the Adam optimizer with a learning rate of $1\times 10^{-2}$.
A.1.4 Dataset Distillation
During dataset distillation, Gaussian points are optimized with the Adam optimizer and a learning rate of $1\times 10^{-2}$. The number of optimization iterations is set to 15,000 unless otherwise specified.
Table 5: Hyperparameters for dataset distillation on CIFAR-10.
Setting GPC num_points syn_steps max_start_epoch expert_epochs lr_lr lr_init batch_syn zca
IPC=1 30 22 60 30 2 1e-5 1e-2 300 True
IPC=10 160 42 60 30 2 1e-5 1e-2 380 True
IPC=50 250 136 60 30 2 1e-5 1e-2 360 True
Table 6: Hyperparameters for dataset distillation on CIFAR-100.
Setting GPC num_points syn_steps max_start_epoch expert_epochs lr_lr lr_init batch_syn zca
IPC=1 30 22 60 30 2 1e-5 1e-2 512 True
IPC=10 80 85 60 30 2 1e-5 1e-2 512 True
IPC=50 400 85 60 30 2 1e-5 1e-2 640 True
Table 7: Hyperparameters for dataset distillation on ImageNet-subset TM).
Setting GPC num_points syn_steps max_start_epoch expert_epochs lr_lr lr_init batch_syn zca
IPC=1 64 170 20 40 2 1e-6 1e-2 150 False
IPC=10 640 170 20 40 2 1e-5 1e-2 160 False
Table 8: Hyperparameters for dataset distillation on ImageNet-subset (DM).
Setting GPC num_points batch_real batch_syn zca Iteration
IPC=1 200 54 1024 2000 False 20000
Table 9: Hyperparameters for dataset distillation on ImageNet-subset (DC).
Setting GPC num_points batch_real batch_syn zca Iteration
IPC=1 200 54 720 2000 False 5000
A.2 More Experimental Results
As shown in Table 11, our method GSDD consistently achieves the best performance across both CIFAR-10 and CIFAR-100 under all IPC settings. Notably, GSDD surpasses all baselines by a clear margin in the low-data regime (IPC=1), achieving 67.6% on CIFAR-10 and 43.0% on CIFAR-100. In high-data settings (IPC=50), GSDD maintains its superiority, reaching 77.7% and 53.1% respectively. These results demonstrate the strong generalization ability and scalability of our primitive-based distillation framework.
We evaluate the generalization ability of distilled data across different architectures, including AlexNet, VGG11, ResNet18, and ViT. The network implementations follow those provided in Cazenavette et al. (2023). For training, the learning rate is set to $1\times 10^{-3}$ for AlexNet, $5\times 10^{-5}$ for ViT, and $1\times 10^{-2}$ for the remaining architectures. All models are optimized using Adam with a cosine annealing learning rate schedule. We report the results averaged over three independent runs, as shown in Figure 10.
Table 10: Cross-architecture results, rounded to one decimal place.
Model Nette Woof Fruit Yellow Meow Squawk
[…]
Mean 58.1 34.6 39.9 53.6 34.0 58.0
Table 11: Results on GSDD on CIFAR-10/100
Method CIFAR10 CIFAR100
IPC=1 IPC=10 IPC=50 IPC=1 IPC=10 IPC=50
Input sized TM 46.3±0.8 65.3±0.7 71.6±0.2 24.3±0.3 40.1±0.4 47.7±0.2
FRePo 46.8±0.7 65.5±0.4 71.7±0.2 28.7±0.1 42.5±0.2 44.3±0.2
Static IDC 50.0±0.4 67.5±0.5 74.5±0.2
FreD 60.6±0.8 70.3±0.3 75.8±0.1 34.6±0.4 42.7±0.2 47.8±0.1
Parameterized HaBa 48.3±0.8 69.9±0.4 74.0±0.2 47.0±0.2
RTP 66.4±0.4 71.2±0.4 73.6±0.5 34.0±0.4 42.9±0.7
HMN 65.7±0.3 73.7±0.1 76.9±0.2 36.3±0.2 45.4±0.2 48.5±0.2
SPEED 63.2±0.1 73.5±0.2 77.7±0.4 40.4±0.4 45.9±0.3 49.1±0.2
NSD 68.5±0.8 73.4±0.2 75.2±0.6 36.5±0.3 46.1±0.2
Function DDiF 66.5±0.4 74.0±0.4 77.5±0.3 42.1±0.2 46.0±0.2 49.9±0.2
Primitive GSDD 67.6±0.4 75.5±0.3 77.7±0.5 43.0±0.1 47.4±0.3 53.1±0.2
A.3 Parallelization Strategy for Batched Rendering
To scale the rendering pipeline from a single image to an entire batch of images while maximizing GPU throughput, we designed a global ID system coupled with corresponding data structure management. The core idea is to consolidate the rendering of multiple independent images into a single, massive computational task.
Core Challenge and the Global ID System
The parallelization strategy for batched rendering evolves the original single-image pipeline’s local indexing system into a global, batch-aware framework. The original approach operated in a local context where screen tiles were indexed from 0 to M-1 for a single image, and data structures pertained only to that instance.
To scale this process across a batch of B images, we introduced a Globally Unique ID (GUID) system. This system creates a single, contiguous address space for all tiles across the entire batch. A local tile j from image i is re-indexed into a global tile id using the linear transformation:
$\text{global\_tile\_id}=i\times M+j$ (12)
This re-indexing prompted adaptations to our key CUDA kernels. The intersection mapping kernel was modified to calculate this global tile id for each generated intersection record, enabling the subsequent sorting operation to correctly group tasks on a global, batch-wide level. Correspondingly, the rasterization kernel is now launched with a 1D grid of B×M thread blocks, where each block’s index directly corresponds to a global tile id. Inside the kernel, this global ID is decomposed back into its constituent image id and local tile id. This allows each thread block to precisely identify which tile of which image it is responsible for, ensuring it writes the final pixel color to the correct memory offset within the batch output tensor.
To support this parallelization strategy, we further employ specific data structures at the Python (PyTorch) level to manage and transfer data efficiently.
1D Flattening and Contiguous Memory Layout
We avoid using Python lists or non-contiguous tensors to store the parameters of different images. Instead, all Gaussian parameters (e.g., means, features, and Cholesky components) across the batch are concatenated and stored in a single large contiguous PyTorch tensor. For a batch of $B$ images with $N$ points each, the means tensor has a shape of $(B\times N,2)$, rather than being represented as a list of $B$ tensors of shape $(N,2)$. This design ensures memory contiguity and eliminates the performance overhead caused by frequent concatenation.
By combining the global ID system at the CUDA kernel level with a contiguous memory layout at the framework level, we have constructed an end-to-end rendering architecture that achieves high throughput for large-scale batched rendering tasks.
A.4 Visualization of Antialias And Gaussian Escape
Figure 6: Comparison between aliasing and anti-aliasing. (a) shows the schematic illustration, while (b) demonstrates the effect on CIFAR-10 dataset visualization. The anti-aliasing strategy leads to smoother and clearer patterns.
(a) CIFAR-10
(b) Imagenette
Figure 7: Visualization of Gaussian point distributions without boundary constraints. We plot the center coordinates of all Gaussian points after training and observe clear escaping phenomena on both CIFAR-10 and Imagenette, which reduces the representational efficiency of Gaussian points.
Figure 6 visualizes the test images without and with anti-aliasing, along with the actual renderings on CIFAR-10. It is evident that when the Gaussian points are relatively elongated, the rendering quality degrades, resulting in pronounced aliasing artifacts and discontinuous stripes, which in turn adversely affect the performance of the distilled dataset.
In Figure 7, we visualize the escaping phenomenon of Gaussian points after training when no boundary constraints are applied. We plot the distributions of all Gaussian point centers and observe that, on both CIFAR-10 and Imagenette datasets, Gaussian points tend to escape beyond the valid range, which in turn undermines their representational efficiency.
A.5 Visualization of GSDD
Figure 8: Visualization of the Gaussian initialization process. The first row shows the actual renderings of Gaussian points at different iterations, while the second row represents each Gaussian as an ellipse encoding its position, shape, and color. During optimization, Gaussian points continuously adjust their positions, shapes, and colors to fit the target image.
We visualize the process of Gaussian initialization in Figure 8. The first row illustrates the actual rendering results of Gaussian points, while the second row depicts the Gaussians as ellipses that encode their positions, shapes, and colors. It can be observed that, during optimization, Gaussian points continuously move, reshape, and adjust their colors to fit the target image.
Figure 9: Visualization of Distilled Images on CIFAR-10
Figure 10: Visualization of Distilled Images on CIFAR-100
Figure 11: Visualization of Distilled Images on ImageNette

### v2

@@
 Experiments show that GSDD achieves state-of-the-art performance on CIFAR-10, CIFAR-100, and ImageNet subsets, while remaining highly efficient encoding and decoding cost.
-Our code is available at https://github.com/j-cyoung/GSDatasetDistillation.
+The code is available at https://github.com/j-cyoung/GSDatasetDistillation.
 1 introduction
-Guided by the principles of scaling laws Kaplan et al.
-(2020), deep learning has advanced along the trajectory of larger models, larger datasets, and longer training, which has significantly pushed the performance boundaries of modern models.
+Guided by the principles of scaling laws kaplan2020scalinglawsneurallanguage, deep learning has advanced along the trajectory of larger models, larger datasets, and longer training, which has significantly pushed the performance boundaries of modern models.
 However, this paradigm also incurs enormous demands on computation, storage, and communication resources, creating a barrier that hinders both broad adoption and sustainable development.
@@
 By adopting dataset distillation, the costs of model training, data storage, and data transmission can be substantially reduced.
-Moreover, it opens up a wide range of applications, including efficient data replay for continual learning Yang et al.
-(2023); Gu et al.
-(2024), critical data transfer in federated learning Huang et al.
-(2024a), and privacy preservation Dong et al.
-(2022).
+Moreover, it opens up a wide range of applications, including efficient data replay for continual learning yang2023an; guSummarizingStreamData2024, critical data transfer in federated learning huang2024overcoming, and privacy preservation dongPrivacyFreeHow2022.
 Dataset distillation optimizes the synthetic data itself, with the objective of minimizing the training-aware information gap between the synthetic and original datasets.
 Broadly, dataset distillation can be decomposed into two complementary modules: distillation algorithms and data parameterization.
-The former defines how training-aware key information, such as training gradients Zhao et al.
-(2021), model parameters Cazenavette et al.
-(2022), or distributional statistics Zhao & Bilen (2023), is extracted, while the latter determines the representation format of the synthetic data.
+The former defines how training-aware key information, such as training gradients zhaoDatasetCondensationGradient2021, model parameters Cazenavette_2022_CVPR, or distributional statistics zhaoDatasetCondensationDistribution2023, is extracted, while the latter determines the representation format of the synthetic data.
 Although extensive research has focused on innovating distillation algorithms, exploration of the representation space has been relatively limited, with most approaches directly optimizing in the pixel space.
@@
 Decoder-based parameterizations, in contrast, can leverage structural priors to enhance image realism and share common information, but they inevitably restrict the optimization space and introduce additional computational and memory overhead.
-More recently, approaches based on implicit neural representations (INRs), such as DDiF Shin et al.
-(2025), have substantially reduced per-image storage and thereby improved the diversity of distilled datasets.
+More recently, approaches based on implicit neural representations (INRs), such as DDiF shin2025distilling, have substantially reduced per-image storage and thereby improved the diversity of distilled datasets.
 However, their inherent per-pixel query decoding mechanism incurs prohibitive computational costs, particularly in high-resolution or large-batch scenarios, creating a severe performance bottleneck.
@@
 Dataset Distillation Algorithms
-Modern deep learning models typically rely on large-scale training datasets, whereas the goal of dataset distillation is to replace the dataset with a much smaller set of high-quality training samplesWang et al.
-(2018).
+Modern deep learning models typically rely on large-scale training datasets, whereas the goal of dataset distillation is to replace the dataset with a much smaller set of high-quality training sampleswangDatasetDistillation2018.
 Unlike data selection, dataset distillation explicitly optimizes the synthetic data to achieve superior performance.
@@
 In dataset distillation, the synthetic dataset is treated as a set of learnable objects, where the pixel values of distilled images $s_{j}$ are regarded as optimization parameters.
-The ultimate goal is to train a model $\theta_{\mathcal{S}}$ on the compact distilled dataset such that its performance closely matches that of a model $\theta_{\mathcal{T}}$ trained on the full dataset Wang et al.
-(2018):
+The ultimate goal is to train a model $\theta_{\mathcal{S}}$ on the compact distilled dataset such that its performance closely matches that of a model $\theta_{\mathcal{T}}$ trained on the full dataset wangDatasetDistillation2018:
 $\mathcal{S}=\arg\min_{\mathcal{S}}\ \big|l(\theta_{\mathcal{S}},\mathcal{D})-l(\theta_{\mathcal{T}},\mathcal{D})\big|,\quad\theta_{\mathcal{S}}=\arg\min_{\theta}l(\theta,\mathcal{S}),\quad\theta_{\mathcal{T}}=\arg\min_{\theta}l(\theta,\mathcal{T}).$
-This optimization objective can be approached via meta-learning, but such methods often consume substantial GPU memory Feng et al.
-(2024) and yield suboptimal performance due to weak supervision signals.
+This optimization objective can be approached via meta-learning, but such methods often consume substantial GPU memory fengEmbarrassinglySimpleDataset2024 and yield suboptimal performance due to weak supervision signals.
 To address these challenges, many alternative distillation objectives have been proposed.
 The core idea is to extract and align critical training-aware information between $\mathcal{S}$ and $\mathcal{T}$.
-Mainstream approaches include trajectory matching(TM) Cazenavette et al.
-(2022); Du et al.
-(2023); Liu et al.
-(2025); Guo et al.
-(2024); Cui et al.
-(2023), which constrains the distance between model parameters trained on the two datasets; distribution matching(DM) Zhao & Bilen (2023); Zhao et al.
-(2023); Liu et al.
-(2023); Wei et al.
-(2024); Wang et al.
-(2025b), which aligns feature distributions; and gradient matching(DC) Zhao et al.
-(2021); Lee et al.
-(2022); DU et al.
-(2023), which constrains the gradients derived from the two datasets.
+Mainstream approaches include trajectory matching(TM) Cazenavette_2022_CVPR; duMinimizingAccumulatedTrajectory2023; liuDatasetDistillationAutomatic2025; guoLosslessDatasetDistillation2024; cuiScalingDatasetDistillation2023, which constrains the distance between model parameters trained on the two datasets; distribution matching(DM) zhaoDatasetCondensationDistribution2023; zhaoImprovedDistributionMatching2023; liuDREAMEfficientDataset2023a; Wei_2024_CVPR; Wang_2025_CVPR, which aligns feature distributions; and gradient matching(DC) zhaoDatasetCondensationGradient2021; leeDatasetCondensationContrastive2022; duuentialSubsetMatching2023, which constrains the gradients derived from the two datasets.
 Despite the drastic reduction in storage requirements, distilled datasets often achieve competitive training performance, and the resulting synthetic images can diverge significantly from the originals.
-Consequently, dataset distillation has found applications in continual learning Yang et al.
-(2023); Gu et al.
-(2024), privacy preservation Dong et al.
-(2022); Chung et al.
-(2024); Zheng et al.
-(2025), federated learning Huang et al.
-(2024a); Jia et al.
-(2025); Yan et al.
-(2025), and neural architecture search Such et al.
-(2020).
+Consequently, dataset distillation has found applications in continual learning yang2023an; guSummarizingStreamData2024, privacy preservation dongPrivacyFreeHow2022; chung2024rethinking; zheng2025improvingnoiseefficiencyprivacypreserving, federated learning huang2024overcoming; UPFL; yan2025fedvck, and neural architecture search such2020generative.
 Parameterization of Dataset Distillation
@@
 Early methods directly optimized RGB pixels, but this dense representation wastes storage by ignoring spatial redundancy.
-To address this, decoder-based approaches Liu et al.
-(2022); Liu & Wang (2023); Cazenavette et al.
-(2023) learn latent codes with a shared decoder, which reduces redundancy but limits flexibility in optimization and adds computation costs.
-Dictionary-based methods Deng & Russakovsky (2022); Wei et al.
-(2023) instead combine basis images with index matrices to support domain adaptation and generalization, though they still store pixel-level bases and are restricted by linear combinations.
-Recent studies have introduced compact low-level representations to reduce image redundancy and increase distilled samples within a fixed storage budget, such as downscaling Kim et al.
-(2022), frequency-domain transforms Shin et al.
-(2023), and implicit neural representations (INRs) Shin et al.
-(2025).
+To address this, decoder-based approaches liuDatasetDistillationFactorization2022; liuMGDDMetaGenerator2023; cazenavetteGeneralizingDatasetDistillation2023 learn latent codes with a shared decoder, which reduces redundancy but limits flexibility in optimization and adds computation costs.
+Dictionary-based methods dengRememberDistillingDatasets2022; weiSparseParameterizationEpitomic2023 instead combine basis images with index matrices to support domain adaptation and generalization, though they still store pixel-level bases and are restricted by linear combinations.
+Recent studies have introduced compact low-level representations to reduce image redundancy and increase distilled samples within a fixed storage budget, such as downscaling kimDatasetCondensationEfficient2022, frequency-domain transforms shinFrequencyDomainBasedDataset2023, and implicit neural representations (INRs) shin2025distilling.
 INR-based methods show strong performance by encoding each image as a neural field with fewer parameters, but their efficiency is limited since decoding requires per-pixel RGB queries, causing heavy overhead for high-resolution and large-batch training.
 Gaussian Splatting
-Gaussian splatting Kerbl et al.
-(2023) represents 3D scenes using Gaussian ellipsoids, offering higher rendering efficiency and explicit geometric control compared to NeRF-based representations Mildenhall et al.
-(2021).
+Gaussian splatting 3dgs represents 3D scenes using Gaussian ellipsoids, offering higher rendering efficiency and explicit geometric control compared to NeRF-based representations NeRF.
 Originally developed for novel view synthesis, it initializes a set of Gaussian primitives within a 3D scene, which are then rendered from multiple viewpoints and optimized to match the corresponding ground-truth images, ultimately enabling arbitrary-view synthesis.
-In recent years, Gaussian splatting has been extended to a wide range of domains, including digital humans Qian et al.
-(2024); Zielonka et al.
-(2025), autonomous driving Zhou et al.
-(2024); Huang et al.
-(2024b), dynamic scene modeling Wu et al.
-(2024); Huang et al.
-(2024c), and 3D editing Chen et al.
-(2024).
-More recently, Gaussian representations have also been applied to 2D vision tasks such as image super-resolution Hu et al.
-(2024); Chen et al.
-(2025), image representation Zhu et al.
-(2025); Weiss & Bradley (2024); Zhang et al.
-(2025), and video representation Wang et al.
-(2025a), where Gaussian primitives are used to efficiently encode high-fideliy image of videos with fine-grained textures.
+In recent years, Gaussian splatting has been extended to a wide range of domains, including digital humans qian20243dgs; zielonka2025drivable, autonomous driving zhou2024drivinggaussian; huang2024textit, dynamic scene modeling wu20244d; huang2024sc, and 3D editing chen2024gaussianeditor.
+More recently, Gaussian representations have also been applied to 2D vision tasks such as image super-resolution hu2024gaussiansrhighfidelity2d; chen2025generalizedefficient2dgaussian, image representation zhu2025largeimagesgaussianshighquality; weiss2024gaussianbillboardsexpressive2d; gaussianimage, and video representation wang2025gsvcefficientvideorepresentation, where Gaussian primitives are used to efficiently encode high-fidelity image of videos with fine-grained textures.
 This work is the first to introduce Gaussian Splatting as a parameterization for dataset distillation, which leverages Gaussian representations as a simple yet efficient parameterization for distilled datasets.
@@
 2.2 Motivation
-Gaussian representations enable faster decoding and rendering, especially compared to the state-of-the-art DDiF Shin et al.
-(2025) method.
-DDiF relies on implicit neural representations using a SIREN network Sitzmann et al.
-(2020) to model distilled images, where each pixel’s RGB value must be queried individually, as illustrated in Figure 1a.
+Gaussian representations enable faster decoding and rendering, especially compared to the state-of-the-art DDiF shin2025distilling method.
+DDiF relies on implicit neural representations using a SIREN network NEURIPS2020_53c04118 to model distilled images, where each pixel’s RGB value must be queried individually, as illustrated in Figure 1a.
 This approach incurs substantial time and memory overhead as the number and resolution of distilled images increase.
@@
 $\mathbf{c}(x,y)=\sum_{k=1}^{M}\alpha_{k}\cdot g_{k}(x,y;u_{k},v_{k},\Sigma_{k})\cdot\mathbf{c}_{k}.$ (4)
-Thus, each Gaussian component can be fully described by a 9-dimensional vector $p_{k}=(u_{k},v_{k},l_{11}^{k},l_{21}^{k},l_{22}^{k},\mathbf{c}_{k},\alpha_{k})\in\mathbb{R}^{9}.$ Accordingly, the parameter set of an entire distilled image $s_{j}$ is given by $\mathbf{p}^{j}=\{p_{k}^{j}\}_{k=1}^{M}$, and the parameterization of the whole distilled dataset $\mathcal{S}$ is $\mathcal{P}_{\mathcal{S}}=\{\mathbf{p}^{j}\}_{j=1}^{N_{\mathcal{S}}}.$ Based on this parameterization, the rendering process of a distilled image can be expressed as a differentiable synthesis function
+Thus, each Gaussian component can be fully described by a 9-dimensional vector $p_{k}=(u_{k},v_{k},l_{11}^{k},l_{21}^{k},l_{22}^{k},\mathbf{c}_{k},\alpha_{k})\in\mathbb{R}^{9}.$ Accordingly, the parameter set of an entire distilled image $s_{j}$ is given by $\mathbf{p}^{j}=\{p_{k}^{j}\}_{k=1}^{M}$, and the parameterization of the whole distilled dataset $\mathcal{S}$ is $\mathcal{P}_{\mathcal{S}}=\{\mathbf{p}^{j}\}_{j=1}^{N_{\mathcal{S}}}.$ Based on this parameterization, $R$ can be seen as a rendering function, which maps the parameters $\mathbf{p}^{j}$ onto a predefined pixel coordinate grid $\mathcal{C}$, and is implemented as the cuda operator,
 $s_{j}=R(\mathbf{p}^{j};\mathcal{C}),$ (5)
@@
 Specifically, all Gaussian primitives across the entire distilled dataset are represented as a single contiguous 1D vector, and a globally unique ID (GUID) system is used to assign threads that rasterize multiple images in parallel.
-Details are provided in Appendix A.3, and code will be open-sourced.
+Details are provided in Appendix C, and code will be open-sourced.
 Some Gaussians may become highly anisotropic during optimization to capture high-frequency features such as edges, which necessitates anti-aliasing strategies.
@@
 4 experiments
-Table 1: Results on ImageNet-Subset, TM
+Table 1: Test accuracy on ImageNet subsets ($128\times 128$) with trajectory matching (TM).
 Subset Nette Woof Fruit Yellow Meow Squawk
@@
 Primitive GSDD 76.4 46.4 51.2 72.4 53.8 73.2
-Table 2: Results with DC/DM
+Table 2: Test accuracy on ImageNet subsets ($128\times 128$) with gradient matching (DC) and distribution matching (DM).
 DC ImageNet Subset ($\times128$) Avg
@@
 Datasets and Baselines
-We evaluate our method on standard datasets: CIFAR-10, CIFAR-100 (at $32\times 32$ resolution), and six ImageNet subsets (at $128\times 128$ resolution).
+We evaluate our method on standard datasets: CIFAR-10, CIFAR-100 (at $32\times 32$ resolution), six ImageNet subsets (at $128\times 128$ resolution and $256\times 256$ resolution), and ImageNet-1K (at $224\times 224$ resolution).
 The general pipeline of dataset distillation involves two stages: (1) generating a synthetic dataset, and (2) training a model from scratch on the synthetic data.
 The performance of the trained model on the real test set is used to assess the quality of the distilled dataset.
-We include a wide range of recent dataset distillation baselines, including TM Cazenavette et al.
-(2022), FRePo Zhou et al.
-(2022), IDC Kim et al.
-(2022), FreD Shin et al.
-(2023), HaBa Liu et al.
-(2022), RTP Deng & Russakovsky (2022), HMN Liu et al.
-(2022), SPEED Wei et al.
-(2023), NSD Yang et al.
-(2025), GLaD Cazenavette et al.
-(2023), H-GLaD Zhong et al.
-(2025), LD3M Moser et al.
-(2024) and DDiF Shin et al.
-(2025).
+We include a wide range of recent dataset distillation baselines, including TM Cazenavette_2022_CVPR, FRePo zhouDatasetDistillationUsing2022, IDC kimDatasetCondensationEfficient2022, FreD shinFrequencyDomainBasedDataset2023, HaBa liuDatasetDistillationFactorization2022, RTP dengRememberDistillingDatasets2022, HMN liuDatasetDistillationFactorization2022, SPEED weiSparseParameterizationEpitomic2023, NSD NSD, GLaD cazenavetteGeneralizingDatasetDistillation2023, H-GLaD Zhong_2025_CVPR, LD3M moser2024latentdatasetdistillationdiffusion and DDiF shin2025distilling.
 Experimental Details
-We use TM Cazenavette et al.
-(2022), DM Zhao & Bilen (2023), and DC Zhao et al.
-(2021) as the underlying distillation algorithms in our experiments.
+We use TM Cazenavette_2022_CVPR, DM zhaoDatasetCondensationDistribution2023, and DC zhaoDatasetCondensationGradient2021 as the underlying distillation algorithms in our experiments.
 All parameters of Gaussians are trained using the Adam optimizer, and the learning rate for Gaussian parameters is uniformly set to 0.001.
@@
 To initialize the Gaussian representation, we randomly sample real images from the original dataset and fit them with Gaussians before entering the standard distillation loop.
-We provide complete implementation details and hyperparameters in the Appendix A.1.
-Table 3: Ablation
+We provide complete implementation details and hyperparameters in the Appendix A.
+Storage Budget and Evaluation Protocol
+We evaluate performance under varying storage budgets, specified in IPC (images per class).
+Since our method utilizes multiple low-storage Gaussian images, we additionally report GPC (Gaussian Images per Class) in Appendix A.
+The number of Gaussians per image is computed as $\text{pts}=\frac{\text{res}\times\text{res}\times 3\times\text{IPC}\times 2}{\text{GPC}\times 9}$, where the factor of 2 accounts for the use of bf16, and the denominator 9 reflects the number of parameters per Gaussian.
+Performance Comparison
+Through the integration of GSDD, our method consistently outperforms the pixel-based TM baseline and other dataset parameterization methods on six ImageNet subsets with IPC=1, as shown in Table 2.
+The results on the higher-resolution ImageNet subset (256×256) are provided in Table 4, where our method remains superior at higher resolutions.
+In addition, we also report experiments conducted at a lower resolution on CIFAR-10/100, with the results presented in Appendix Table 13.
+When compared to the state-of-the-art method DDiF, our approach achieves consistent improvements, demonstrating highly competitive performance.
+These results highlight the effectiveness of GSDD in enhancing the quality of distilled datasets.
+We further report results on ImageNet-1K($224\times 224$) in Table 17 in Appendix B.4.
+Universality to Matching Objectives
+GSDD is designed to be compatible with a wide range of dataset distillation algorithms and can be directly integrated with them.
+We evaluate GSDD under two widely adopted loss paradigms: gradient matching (DC) and distribution matching (DM).
+Experimental results, as shown in Table 2, confirm that GSDD robustly enhances performance across different distillation objectives.
+Cross-Architecture Performance
+An essential property of dataset distillation is robustness across model architectures.
+The distilled data should retain high performance even when the evaluation model differs from the one used during distillation.
+To assess this, we synthesize data using a ConvNet and evaluate it on a variety of downstream architectures, including ResNet, VGG, AlexNet and ViT.
+The results, summarized in Table 4 and Table 14, demonstrate that the increased diversity introduced by GSDD leads to substantial improvements in cross-architecture generalization.
+Figure 5: Performance comparison between GSDD and DDiF under the same per-image storage budget (abbreviated as st (floats)).
+Top row: Forward and forward+backward execution time and memory usage across varying image resolutions (with fixed batch size = 32).
+Bottom row: Same metrics across varying batch sizes (with fixed resolution = 128).
+GSDD consistently achieves lower latency and memory consumption, especially under high-resolution and large-batch scenarios.
+Table 3: Results on ImageNet-subsets with Distribution Matching (DM) for $256\times 256$
+Method Nette Woof Fruit Yellow Meow Squawk
+Vanilla 32.1 20.0 19.5 33.4 21.2 27.6
+IDC 53.7 30.2 33.1 52.2 34.6 47.0
+FreD 54.2 31.2 32.5 49.1 34.0 43.1
+LatentDD 56.1 28.0 30.7 36.3 47.1
+DDiF 67.8 39.6 43.2 63.1 44.8 67.0
+GSDD 70.0 42.6 51.2 67.4 46.4 70.4
+Table 4: Cross-Architecture Performance
+Subset Nette Woof Fruit Yellow Meow Squawk
+TM 22.0 14.8 17.1 22.3 16.2 25.5
+IDC 27.9 19.5 23.9 28.0 19.8 29.9
+FreD 36.2 23.7 23.6 31.2 19.1 37.4
+GLaD 30.4 17.1 21.1 19.6 28.2
+H-GLaD 30.8 17.4 21.5 20.1 28.8
+LD3M 32.0 19.9 21.4 22.1 30.4
+DDiF 59.3 34.1 39.3 51.1 33.8 54.0
+GSDD 58.1 34.6 39.9 53.6 34.0 58.0
+4.2 Benchmark Rendering Efficiency
+To evaluate the efficiency and representational capacity of GSDD, we conduct an image-fitting experiment comparing several recent parameterization methods with the same storage budget, including FreD shinFrequencyDomainBasedDataset2023, SPEED weiSparseParameterizationEpitomic2023 and DDiF shin2025distilling.
+Specifically, we use the ImageNette dataset and fit the first 100 images from each class (1,000 images in total) at a resolution of 128.
+The optimization is performed for 1,000 steps, and for each method we perform a grid search over learning rates and optimizers to ensure a fair comparison.
+As shown in Table 5, GSDD shows great performance across representational quality (PSNR), GPU memory usage, and runtime efficiency.
+High fidelity and low computational overhead of GSDD enables fast synthesis of more diverse (Higher GPC) distilled datasets, which in turn leads to improved downstream performance.
+We further verify in Section 4.3 that larger GPC values with fixed IPC generally correspond to stronger distilled data performance.
+Table 5: Comparison of Parameterization Methods on Image Fitting Efficiency and Representational Quality
+FreD SPEED DDiF GSDD
+PSNR(dB)↑ 15.69 17.77 23.23 23.22
+VRAM(MB)↓ 280.4 298.5 1078.8 150.2
+Time(s)↓ 15.5 69.4 1088.4 24.5
+We further conduct a comprehensive quantitative comparison with the state-of-the-art DDiF method in terms of both decoding and training performance.
+As shown in Figure 5, we measure the inference time and memory consumption of both methods across varying per-image storage budgets, image resolutions, and batch sizes.
+The results demonstrate that GSDD consistently outperforms DDiF in both inference/training speed and memory efficiency.
+This advantage becomes especially pronounced when handling high-resolution inputs or large batch sizes, where GSDD achieves several-fold, or even an order-of-magnitude, reductions in both computation time and memory usage.
+These findings indicate that our method not only maintains strong image representation capabilities for distillation but also offers superior rendering speed and significantly lower memory footprint, thereby substantially improving the scalability of dataset distillation methods in practical deployment scenarios.
+4.3 Ablation Study
+Table 6: Ablation Study on Individual Components
 ImageNette ACC
@@
 all 75.5
-Storage Budget and Evaluation Protocol
-We evaluate performance under varying storage budgets, specified in IPC (images per class).
-Since our method often utilizes multiple low-storage Gaussian images, we additionally report GPC (Gaussian Images per Class) in Appendix A.1.
-The number of Gaussians per image is computed as $\text{pts}=\frac{\text{res}\times\text{res}\times 3\times\text{IPC}\times 2}{\text{GPC}\times 9}$, where the factor of 2 accounts for our use of bf16 to store each Gaussian parameter, and the denominator 9 reflects the total number of parameters per Gaussian.
-Performance Comparison
-Through the integration of GSDD, our method consistently outperforms the pixel-based TM baseline and other dataset parameterization methods on six ImageNet subsets with IPC=1, as shown in Table 2.
-Extensive results on CIFAR-10 and CIFAR-100 with IPC values of 1, 10, and 50 are provided in Appendix A.2.
-When compared to the state-of-the-art method DDiF, our approach achieves consistent improvements, demonstrating highly competitive performance.
-These results highlight the effectiveness of GSDD in enhancing the quality of distilled datasets.
-Table 4: Cross Architecture Results
-Subset Nette Woof Fruit Yellow Meow Squawk
-TM 22.0 14.8 17.1 22.3 16.2 25.5
-IDC 27.9 19.5 23.9 28.0 19.8 29.9
-FreD 36.2 23.7 23.6 31.2 19.1 37.4
-GLaD 30.4 17.1 21.1 19.6 28.2
-H-GLaD 30.8 17.4 21.5 20.1 28.8
-LD3M 32.0 19.9 21.4 22.1 30.4
-DDiF 59.3 34.1 39.3 51.1 33.8 54.0
-GSDD 58.1 34.6 39.9 53.6 34.0 58.0
-Universality to Matching Objectives
-GSDD is designed to be compatible with a wide range of dataset distillation algorithms and can be directly integrated with them.
-We evaluate GSDD under two widely adopted loss paradigms: gradient matching(DC) and distribution matching(DM).
-Experimental results, as shown in Table 2, confirm that GSDD robustly enhances performance across different distillation objectives.
-Cross-Architecture Performance
-An essential property of dataset distillation is robustness across model architectures.
-The distilled data should retain high performance even when the evaluation model differs from the one used during distillation.
-To assess this, we synthesize data using a ConvNet and evaluate it on a variety of downstream architectures, including ResNet, VGG, and others.
-The results, summarized in Table 4, demonstrate that the increased diversity introduced by GSDD leads to substantial improvements in cross-architecture generalization.
-4.2 More Analysis
-Figure 5: Performance comparison between GSDD and DDiF under the same per-image storage budget (abbreviated as st (floats)).
-Top row: Forward and forward+backward execution time and memory usage across varying image resolutions (with fixed batch size = 32).
-Bottom row: Same metrics across varying batch sizes (with fixed resolution = 128).
-GSDD consistently achieves lower latency and memory consumption, especially under high-resolution and large-batch scenarios.
 To investigate the contribution of each component to model performance, we conduct a series of ablation studies, with results summarized in Table 6.
 The experiments show that opacity modeling and bf16 mixed-precision training are two major factors, as removing either leads to a notable performance drop.
+The boundary entry in the table refers to the result obtained when both hard clipping and loss regularization are removed.
 The impact of anti-aliasing exhibits a dependence on the dataset and the complexity of the representation.
@@
 In contrast, for datasets like ImageNette with denser Gaussians, aliasing artifacts are largely mitigated, and the performance gains from anti-aliasing become relatively minor.
-To comprehensively evaluate the efficiency of our approach, we conduct a quantitative comparison with the state-of-the-art DDiF method in terms of both decoding and training performance.
-As shown in Figure 5, we measure the inference time and memory consumption of both methods across varying per-image storage budgets, image resolutions, and batch sizes.
-The results demonstrate that GSDD consistently outperforms DDiF in both inference/training speed and memory efficiency.
-This advantage becomes especially pronounced when handling high-resolution inputs or large batch sizes, where GSDD achieves several-fold, or even an order-of-magnitude, reductions in both computation time and memory usage.
-These findings indicate that our method not only maintains strong image representation capabilities for distillation but also offers superior rendering speed and significantly lower memory footprint, thereby substantially improving the scalability of dataset distillation methods in practical deployment scenarios.
+We further examine the effect of GPC through an ablation study, as shown in Figure 6.
+When using the memory-efficient DM algorithm or when distilling under small storage budgets, performance increases with GPC at first but eventually declines.
+This is because small GPC values lead to insufficient diversity, whereas excessively large GPC values allocate too few Gaussians per image, weakening each image’s representational capacity.
+In more memory-consuming settings, the optimal turning point of GPC becomes difficult to reach, and performance tends to improve monotonically as GPC increases.
+Figure 6 also shows that the performance of the initial dataset serves as a useful reference for estimating a reasonable GPC range.
+Figure 6: Effect of GPC (Gaussian Images Per Class) on distilled dataset performance across different datasets, storage budgets, and distillation algorithms (TM Cazenavette_2022_CVPR and DM zhaoDatasetCondensationDistribution2023).
+Initial Accuracy denotes the performance of initialized gaussian images (initialized on sampled real images).
+Distilled Accuracy denotes the performance after distillation.
 5 conclusion
@@
 Future work may focus on scaling to larger datasets, extending Gaussian representations to video modalities, and exploring dynamic density control of Gaussian primitives to further enhance representation quality.
-Appendix A Appendix
-A.1 Experimental Details
-A.1.1 Datasets
+Appendix A Experimental Details
+A.1 Datasets
 We conduct experiments on eight datasets in total, including CIFAR-10, CIFAR-100, and six ImageNet subsets: imagenette, imagewoof, imagefruit, imagemeow, imagesquawk, and imageyellow.
@@
 Each ImageNet subset contains images of resolution $128\times 128$ from 10 classes, with more than 10,000 images per subset.
-A.1.2 Network Architectures
+A.2 Network Architectures
 We adopt ConvNet architectures as the backbone networks.
@@
 Specifically, we use ConvNetD3 for CIFAR-10 and CIFAR-100, and ConvNetD5 for the ImageNet subsets.
-A.1.3 Image Initialization
+A.3 Image Initialization
 For the initialization of synthetic images, all experiments adopt the mean squared error (MSE) loss and use the Adam optimizer with a learning rate of $1\times 10^{-2}$.
-A.1.4 Dataset Distillation
+A.4 Dataset Distillation
 During dataset distillation, Gaussian points are optimized with the Adam optimizer and a learning rate of $1\times 10^{-2}$.
 The number of optimization iterations is set to 15,000 unless otherwise specified.
-Table 5: Hyperparameters for dataset distillation on CIFAR-10.
+Table 7: Hyperparameters for dataset distillation on CIFAR-10 $32\times 32$.
 Setting GPC num_points syn_steps max_start_epoch expert_epochs lr_lr lr_init batch_syn zca
@@
 IPC=50 250 136 60 30 2 1e-5 1e-2 360 True
-Table 6: Hyperparameters for dataset distillation on CIFAR-100.
+Table 8: Hyperparameters for dataset distillation on CIFAR-100 $32\times 32$.
 Setting GPC num_points syn_steps max_start_epoch expert_epochs lr_lr lr_init batch_syn zca
@@
 IPC=50 400 85 60 30 2 1e-5 1e-2 640 True
-Table 7: Hyperparameters for dataset distillation on ImageNet-subset TM).
+Table 9: Hyperparameters for dataset distillation on ImageNet-subset TM $128\times 128$.
 Setting GPC num_points syn_steps max_start_epoch expert_epochs lr_lr lr_init batch_syn zca
@@
 IPC=10 640 170 20 40 2 1e-5 1e-2 160 False
-Table 8: Hyperparameters for dataset distillation on ImageNet-subset (DM).
+Table 10: Hyperparameters for dataset distillation on ImageNet-subset $128\times 128$ (DM).
 Setting GPC num_points batch_real batch_syn zca Iteration
 IPC=1 200 54 1024 2000 False 20000
-Table 9: Hyperparameters for dataset distillation on ImageNet-subset (DC).
+Table 11: Hyperparameters for dataset distillation on ImageNet-subset $256\times 256$ (DM).
+Setting GPC num_points batch_real batch_syn zca Iteration
+IPC=1 120 364 512 512 False 20000
+Table 12: Hyperparameters for dataset distillation on ImageNet-subset $128\times 128$ (DC).
 Setting GPC num_points batch_real batch_syn zca Iteration
 IPC=1 200 54 720 2000 False 5000
-A.2 More Experimental Results
+A.5 Performance Benchmarking
+In Table 5, we evaluate the image-fitting performance of FreD shinFrequencyDomainBasedDataset2023, SPEED weiSparseParameterizationEpitomic2023, DDiF shin2025distilling, and GSDD.
+We select 100 images from each class of the ImageNette dataset, resulting in a total of 1,000 images, all at a resolution of 128.
+Each method is constrained to the same storage budget, meaning that, on average, each image is represented using 963 floating-point parameters.
+For SPEED, we follow the parameter-counting formula provided in its original paper,
+$\text{\#Params}=DK+1.5NHk+R(3D^{2}2+7D)+L(D+1),$
+and, given the storage constraint, we solve for all feasible hyperparameter combinations.
+The final configuration used in our experiments is (D=32, N=32, k=135, R=2, L=64).
+We attempted to evaluate HaBa liuDatasetDistillationFactorization2022 as well, but its model structure cannot be solved inversely under the resolution of 128 and the storage budget of 963 parameters.
+FreD, DDiF, and GSDD exhibit clear advantages in terms of estimating and meeting the storage budget.
+For each method, we search over Adam and SGD optimizers and learning rates spaced exponentially by a factor of 10, and we report the highest PSNR achieved.
+All models are trained for 1,000 iterations using MSE loss.
+Each experiment is repeated three times with random initialization, and the average result is reported.
+The results demonstrate that our method achieves strong representational capability and computational efficiency.
+Appendix B More Experimental Results
+B.1 Performance Comparison on Low-Dimensional Datasets
 As shown in Table 13, our method GSDD consistently achieves the best performance across both CIFAR-10 and CIFAR-100 under all IPC settings.
@@
 These results demonstrate the strong generalization ability and scalability of our primitive-based distillation framework.
-We evaluate the generalization ability of distilled data across different architectures, including AlexNet, VGG11, ResNet18, and ViT.
-The network implementations follow those provided in Cazenavette et al.
-(2023).
-For training, the learning rate is set to $1\times 10^{-3}$ for AlexNet, $5\times 10^{-5}$ for ViT, and $1\times 10^{-2}$ for the remaining architectures.
-All models are optimized using Adam with a cosine annealing learning rate schedule.
-We report the results averaged over three independent runs, as shown in Figure 10.
-Table 10: Cross-architecture results, rounded to one decimal place.
-Model Nette Woof Fruit Yellow Meow Squawk
-[…]
-[…]
-[…]
-[…]
-Mean 58.1 34.6 39.9 53.6 34.0 58.0
-Table 11: Results on GSDD on CIFAR-10/100
+Table 13: Classification accuracy on CIFAR-10 and CIFAR-100 under different IPC settings.
 Method CIFAR10 CIFAR100
@@
 Primitive GSDD 67.6±0.4 75.5±0.3 77.7±0.5 43.0±0.1 47.4±0.3 53.1±0.2
-A.3 Parallelization Strategy for Batched Rendering
+B.2 Detailed Cross-architecture Results
+We evaluate the generalization ability of distilled data across different architectures, including AlexNet, VGG11, ResNet18, and ViT.
+The network implementations follow those provided in cazenavetteGeneralizingDatasetDistillation2023.
+For training, the learning rate is set to $1\times 10^{-3}$ for AlexNet, $5\times 10^{-5}$ for ViT, and $1\times 10^{-2}$ for the remaining architectures.
+All models are optimized using Adam with a cosine annealing learning rate schedule.
+We report the results averaged over three independent runs, as shown in Figure 14.
+Table 14: Detailed cross-architecture performance comparison.
+Test Net Method Nette Woof Fruit Yellow Meow Squawk
+AlexNet TM 13.2 ± 0.6 10.0 ± 0.0 10.0 ± 0.0 11.0 ± 0.2 9.8 ± 0.0
+IDC 17.4 ± 0.9 16.5 ± 0.7 17.9 ± 0.7 20.6 ± 0.9 16.8 ± 0.5 20.7 ± 1.0
+FreD 35.7 ± 0.4 23.9 ± 0.7 15.8 ± 0.7 19.8 ± 1.2 14.4 ± 0.5 36.3 ± 0.3
+DDiF 60.7 ± 2.3 36.4 ± 2.3 41.8 ± 0.6 56.2 ± 0.8 40.3 ± 1.9 60.5 ± 0.4
+GSDD 63.1 ± 1.3 33.6 ± 0.9 41.9 ± 2.0 53.5 ± 0.9 38.1 ± 0.8 60.1 ± 0.2
+VGG11 TM 17.4 ± 2.1 12.6 ± 1.8 11.8 ± 1.0 16.9 ± 1.1 13.8 ± 1.3
+IDC 19.6 ± 1.5 16.0 ± 2.1 13.8 ± 1.3 16.8 ± 3.5 13.1 ± 2.0 19.1 ± 1.2
+FreD 21.8 ± 2.9 17.1 1.7 12.6 ± 2.6 18.2 ± 1.1 13.2 ± 1.9 18.6 ± 2.3
+DDiF 53.6 ± 1.5 29.9 ± 1.9 33.8 ± 1.9 44.2 ± 1.7 32.0 ± 1.8 37.9 ± 1.5
+GSDD 56.8 ± 1.6 32.7 ± 1.4 34.1 ± 3.0 57.3 ± 4.5 31.3 ± 1.8 55.9 ± 3.4
+ResNet18 TM 34.9 ± 2.3 20.7 ± 1.0 23.1 ± 1.5 43.4 ± 1.1 22.8 ± 2.2
+IDC 43.6 ± 1.3 23.2 ± 0.8 32.9 ± 2.8 44.2 ± 3.5 28.2 ± 0.5 47.8 ± 1.9
+FreD 48.8 ± 1.8 28.4 ± 0.6 34.0 ± 1.9 49.3 ± 1.1 29.0 ± 1.8 50.2 ± 0.8
+DDiF 63.8 ± 1.8 37.5 ± 1.9 42.0 ± 1.9 55.9 ± 1.0 35.8 ± 1.8 62.6 ± 1.5
+GSDD 59.1 ± 3.0 39.5 ± 2.0 42.6 ± 0.8 55.9 ± 1.2 40.9 ± 2.9 62.6 ± 0.6
+ViT TM 22.6 ± 1.1 15.9 ± 0.4 23.3 ± 0.4 18.1 ± 1.3 18.6 ± 0.9
+IDC 31.0 ± 0.6 22.4 ± 0.8 31.1 ± 0.8 30.3 ± 0.6 21.4 ± 0.7 32.2 ± 1.2
+FreD 38.4 ± 0.7 25.4 ± 1.7 31.9 ± 1.4 37.6 ± 2.0 19.7 ± 0.8 44.4 ± 1.0
+DDiF 59.0 ± 0.4 32.8 ± 0.8 39.4 ± 0.8 47.9 ± 0.6 27.0 ± 0.6 54.8 ± 1.1
+GSDD 53.4 ± 0.6 32.6 ± 1.1 41.0 ± 1.2 47.6 ± 0.8 25.8 ± 0.4 53.5 ± 2.1
+B.3 Different Initialization Strategy
+Table 15: Initialization Strategy Ablation Results
+Init Acc.
+IDC real 61.4
+FreD real 66.8
+NSD random 68.6
+DDiF real 72.0
+GSDD random 63.6
+GSDD warmup 69.4
+GSDD real 76.4
+Prior methods often initialize distilled data using real images, which may introduce additional privacy risks.
+We further conduct an ablation study on different initialization strategies, including random initialization.
+However, we find that optimizing Gaussian representations from purely random initialization is inherently unstable: the initial Gaussian points are extremely small and sparsely distributed, creating large gaps that prevent effective gradient propagation.
+To mitigate this issue, we introduce a Solid-Color Warmup strategy.
+Instead of fitting any real images, all Gaussian representations are first optimized to fit a single solid-color image.
+This process reduces point sparsity and adjusts the initial point sizes, after which standard distillation begins.
+As shown in Table 15, even without real-image initialization, our warmup-based strategy achieves strong results and surpasses several methods that depend on real-image initialization.
+B.4 Results on ImageNet-1K
+Table 16: Performance on ImageNet-1K.
+IPC=1 IPC=10
+SRe2L 0.9 21.3
+SRe2L+IDC 13.4 41.5
+SRe2L+GSDD 29.0 46.2
+Table 17: Hyperparameters for each method on ImageNet-1K.
+IPC=1 IPC=10
+SRe2L+IDC k=20 k=5
+SRe2L+GSDD GPC=50 GPC=200
+Enabled by the efficiency of recent dataset-distillation frameworks, such as SRe2L yinSqueezeRecoverRelabel2023 and RDED sunDiversityRealismDistilled2024, dataset parameterization at the ImageNet-1K 5206848 scale has become feasible for the first time.
+In this work, we build upon SRe2L and evaluate our method under both IPC=1 and IPC=10 settings.
+In addition, we reproduce the IDC kimDatasetCondensationEfficient2022 baseline, where each image is spatially downscaled by a factor of $\sqrt{k}$, resulting in $k$-times more synthetic samples.
+For IDC, under IPC=1 we report the best-performing $k$ selected from $k\in\{4,20,50\}$, and under IPC=10 from $k\in\{4,10,20\}$.
+As shown in Table 17, our method improves the performance of SRe2L on ImageNet-1K, particularly in the IPC=1 setting.
+Although the gains diminish at IPC=10, our approach still maintains a clear margin over IDC.
+Appendix C Parallelization Strategy for Batched Rendering
 To scale the rendering pipeline from a single image to an entire batch of images while maximizing GPU throughput, we designed a global ID system coupled with corresponding data structure management.
 The core idea is to consolidate the rendering of multiple independent images into a single, massive computational task.
-Core Challenge and the Global ID System
+C.1 Core Challenge and the Global ID System
 The parallelization strategy for batched rendering evolves the original single-image pipeline’s local indexing system into a global, batch-aware framework.
@@
 To support this parallelization strategy, we further employ specific data structures at the Python (PyTorch) level to manage and transfer data efficiently.
-1D Flattening and Contiguous Memory Layout
+C.2 1D Flattening and Contiguous Memory Layout
 We avoid using Python lists or non-contiguous tensors to store the parameters of different images.
@@
 By combining the global ID system at the CUDA kernel level with a contiguous memory layout at the framework level, we have constructed an end-to-end rendering architecture that achieves high throughput for large-scale batched rendering tasks.
-A.4 Visualization of Antialias And Gaussian Escape
+C.3 Benchmark across different GPUs
+To evaluate the stability of our CUDA-based rasterizer, we conduct systematic benchmarking on five different GPU architectures.
+We use an image-fitting task in which Gaussian representations reconstruct 1000 images sampled from ImageNet (10 classes $\times$ 100 images in total).
+Each experiment is repeated three times, and we average all results.
+We test four storage budgets (21, 170, 682, 2730 Gaussians per image).
+All experiments used PyTorch 2.5.1 and CUDA $\leq$ 12.0.
+Table 18: Runtime, memory usage, and PSNR of our rasterizer benchmarked across five GPU architectures under varying storage budgets.
+Time (s) Memory (GB) PSNR
+GPU Model 21 170 682 2730 21 170 682 2730 21 170 682 2730
+Tesla-V100-SXM2-32GB 23.65 26.11 30.28 42.94 1.66 1.70 1.81 2.28 20.24 25.93 30.65 36.09
+NVIDIA-A100-SXM4-40GB 10.45 12.87 19.09 41.53 1.66 1.70 1.81 2.28 20.24 26.03 31.06 36.08
+NVIDIA-GeForce-RTX-3090 10.33 16.07 25.44 61.38 1.66 1.70 1.81 2.28 20.24 26.03 31.06 36.08
+NVIDIA-GeForce-RTX-4090 7.60 9.38 14.06 31.00 1.66 1.70 1.81 2.28 20.24 26.02 31.03 36.06
+NVIDIA-GeForce-RTX-5090 7.13 6.99 9.54 21.79 1.66 1.70 1.81 2.28 20.23 26.02 31.03 36.06
+As shown in Table 18.
+Different GPUs exhibit different runtimes due to heterogeneous compute capabilities, and RTX5090 is the fastest, while RTX3090 is the slowest.
+Memory usage remained identical across all platforms, and PSNR variation is small across architectures.
+Slight performance drops on V100 are expected due to older hardware, but results on A100/3090/4090/5090 remain highly consistent.
+C.4 Results on Cross-Resolution Generalization
+Gaussian Splatting naturally supports rendering at arbitrary resolutions through supersampling, enabling us to distill data at a low resolution (e.g., 128) while still using it to train models at a higher resolution (e.g., 256).
+This effectively reduces the computation when distilling datasets at large resolutions.
+Following results and setup in DDiF shin2025distilling, we conduct cross-resolution experiments as shown in Table 19.
+Table 19: Test accuracies with different resolutions and networks. Images are first distilled with low resolution(128) and test at higher resolution(256/512).
+test resolution test network method accuracy↑ difference↓ ratio↓
+256 ConvNetD5 Vanilla 31.2 20.2 0.39
+IDC 55.0 6.4 0.10
+SPEED 58.8 8.1 0.12
+FreD 56.4 10.4 0.16
+DDiF 66.3 5.7 0.08
+GSDD 66.4 10.0 0.13
+ConvNetD6 Vanilla 44.0 7.3 0.14
+IDC 55.4 6.0 0.10
+SPEED 62.6 4.3 0.06
+FreD 61.8 5.0 0.07
+DDiF 70.6 1.4 0.02
+GSDD 71.6 4.8 0.06
+512 ConvNetD5 Vanilla 27.4 24.0 0.47
+IDC 39.5 21.9 0.36
+SPEED 45.0 21.9 0.33
+FreD 42.9 23.9 0.36
+DDiF 58.7 13.3 0.18
+GSDD 59.4 17.0 0.22
+ConvNetD6 Vanilla 41.2 10.1 0.20
+IDC 51.5 9.9 0.16
+SPEED 60.1 6.8 0.10
+FreD 56.3 10.5 0.16
+DDiF 69.0 3.0 0.04
+GSDD 67.8 8.6 0.11
+For IDC, SPEED, and FreD, we adopt their best-performing upsampling strategies.
+For DDiF and GSDD, supersampling is applied, allowing lossless upscaling of the distilled images.
+GSDD maintains competitive performance in cross-resolution tasks.
+Although its relative performance drop is larger than DDiF, which is likely due to GSDD’s inherent sparsity, reducing its ability to capture fine-grained features when evaluated at higher resolutions.
+Appendix D Visualization of Antialias And Gaussian Escape
 Figure 7: Comparison between aliasing and anti-aliasing.
@@
 We plot the distributions of all Gaussian point centers and observe that, on both CIFAR-10 and Imagenette datasets, Gaussian points tend to escape beyond the valid range, which in turn undermines their representational efficiency.
-A.5 Visualization of GSDD
+Appendix E Visualization of GSDD
 Figure 9: Visualization of the Gaussian initialization process.